The height of a right triangle is 6 inches longer than the base. The area of the triangle is 108 sq in.
Find the measures of the triangles base and height.

Answer by nerdybill(7384)


The height of a right triangle is 6 inches longer than the base. The area of the triangle is 108 sq in.
Find the measures of the triangles base and height.
.
Area of any triangle is:
(1/2)bh
where
b is base
h is height
.
so, now we have:
Let b = length of base
then
b+6 = length
.
So
(1/2)b(b+6) = 108
b(b+6) = 216
b^2+6b = 216
b^2+6b-216 = 0
(b+18)(b-12) = 0
b = {-18, 12}
We can throw out the negative solution leaving:
b = 12 inches (base)
.
Height:
b+6 = 12+6 = 18 inches